OpenStudy (anonymous):

Determine the De Broglie wavelength for a piece of rock that weighs 842 pounds, falling down at the point where it hits the ground after 200 feet.

OpenStudy (michele_laino):

the speed of the rock when it hits the Surface, is: \[\Large v = \sqrt {2gd} \] where \(g\) is gravity, and \(d=200\) feet the De Broglie wavelength \(\lambda\), is: \[\Large \lambda = \frac{h}{{\frac{{Wv}}{g}}}\] where \(W=842\) pounds is the weight of the rock

OpenStudy (anonymous):

may i ask why you divided the de broglie equation by g, since it normally is \[\lambda =h/mv\] is it always when an object is falling that you add the g?

OpenStudy (anonymous):

also, for the speed of the rock would we just use v=\[\sqrt{2gh}\] or should we take the average and divide by 2 ?

OpenStudy (michele_laino):

first question: I have divided by \(g\) since we have: \(W=mg\), then \(m=W/g\)

OpenStudy (michele_laino):

second question: my result above gives the wavelength at the time when the rock hits the earth surface

OpenStudy (anonymous):

oh ok! thanks a lot :)

OpenStudy (michele_laino):

:)